OpenStudy (anonymous):

Factor x2 − 10x + 16.

OpenStudy (anonymous):

If you mean x^2, I believe the answer is (x-8)(x-2)=0 ...

OpenStudy (anonymous):

The factorization of x2 − 7x + 10 = 0 is (x − 5)(x − 2) = 0 True of false

OpenStudy (anonymous):

That is True

OpenStudy (anonymous):

Solve the quadratic equation 3x2 + 5x + 2 = 0 by factoring.

OpenStudy (anonymous):

This might help http://answers.yahoo.com/question/index?qid=20130404181650AA18ZFo And I need to be going now, good luck with the rest of your questions (:

OpenStudy (anonymous):

Thanks